8. Healthy Practices Hospital purchases $300,000 of testing equipment that has a useful life of 5 years. The hospital also pays $20,000 to install the equipment and make it usable. The hospital also expects that after 5 years, it can sell the equipment for $10,000. Calculate the five years of depreciation using straight line, double declining balance, and sum-of-the-years digits.\\
Answer #1
Cost of the equipment=Acquisition cost+Installation cost=300000+20000=$ 320000
Useful life=5 years
Salvage value=$ 10000
Straight-line:
Depreciation under straight-line method=(Cost-Salvage value)/Useful life
It will be same for all 5 years
Depreciation under straight-line method=(320000-10000)/5=$ 62000
Double declining balance:
Depreciation rate=2*(1/Useful life)=2*(1/5)=2*0.2=0.4=40%
Depreciation for first year=Cost*Depreciation rate
Depreciation for rest of the years=Beginning book value*Depreciation rate
Year 1:
Depreciation=320000*40%=$ 128000
Year 2:
Beginning book value=Cost-Depreciation for year 1=320000-128000=$ 192000
Depreciation=192000*40%=$ 76800
Year 3:
Beginning book value=Beginning book value for year 2-Depreciation for year 2=192000-76800=$ 115200
Depreciation=115200*40%=$ 46080
Year 4:
Beginning book value=Beginning book value for year 3-Depreciation for year 3=115200-46080=$ 69120
Depreciation=69120*40%=$ 27648
Year 5:
Beginning book value=Beginning book value for year 4-Depreciation for year 4=69120-27648=$ 41472
Depreciation=41472*40%=$ 16589
Sum-of the years digits:
Sum-of the years digits=5+4+3+2+1=15
Depreciation=(Cost-Salvage value)*Depreciation rate
Year 1:
Depreciation rate=5/15=0.3333=33.33%
Depreciation=(320000-10000)*33.33%=$ 103323
Year 2:
Depreciation rate=4/15=0.2667=26.67%
Depreciation=(320000-10000)*26.67%=$ 82677
Year 3:
Depreciation rate=3/15=0.2=20%
Depreciation=(320000-10000)*20%=$ 62000
Year 4:
Depreciation rate=2/15=0.1333=13.33%
Depreciation=(320000-10000)*13.33%=$ 41323
Year 5:
Depreciation rate=1/15=0.0667=6.67%
Depreciation=(320000-10000)*6.67%=$ 20677